OpenStudy (anonymous):

find the x-intercept of the function y=x^2 + x - 2

OpenStudy (anonymous):

x intercept means y = 0 x^2 + x -2 = 0 can you solve it ?

OpenStudy (anonymous):

no how do i solve it?

OpenStudy (anonymous):

two options .. you may use : ax^2 + bx + c = 0 \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

in this case a = 1 , b = 1 , c = -2

OpenStudy (anonymous):

another way is factoring it directly

OpenStudy (anonymous):

the square root is above b^2-4ac

OpenStudy (anonymous):

dont know why it looks like its above b^2

OpenStudy (anonymous):

can you do it using the formula i wrote ?

OpenStudy (anonymous):

In this case I would definitely factor. Very quick and easy.

OpenStudy (anonymous):

yes, i am trying to factor it though thats the way my teacher wants it done.

OpenStudy (anonymous):

i would get (x+2)(x-1)?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok so in order to factor ax^2 + bx + c you have to split the middle term 'b' into b1 and b2 those b1 and b2 should satisfy : b1 + b2 = b b1 * b2 = a *c in this case we can do : x^2 + x -2 = 0 see here b =2 so we choose b1 =2 , b2 -1 b1 + b2 = 1 b1 *b2 = -2 then it becomes : x^2 + 2x - x -2 = 0 can you factor it now ?

OpenStudy (anonymous):

So the solutions are: x+2 = 0 x-1 = 0

OpenStudy (anonymous):

oh lol havent seen that you wrote it already! yes (x+2)(x-1) = 0 now you know the solutions (x that satisfy the equation) ?

OpenStudy (anonymous):

so there are two x-intercepts?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

thanks!. so the intercepts would be (-2,0) and (1,0)?

OpenStudy (anonymous):

yes.. very good

OpenStudy (anonymous):

Myep.